OpenStudy (anonymous):

adding and subtracting rational expressions. simplify. ( t^2 + 6t/ t - 1) - (7/ t - 1)

OpenStudy (tkhunny):

Gaa!! Please remember your Order of Operations. You have written this: \((t^{2} + 6 - 1)-\left(\dfrac{7}{t} - 1\right)\). It would help to use parentheses where they are NECESSARY. (t^2 + 6t)/ (t - 1) - 7/( t - 1) This means \(\dfrac{t^{2} + 6t}{t-1} - \dfrac{7}{t-1}\) As such, they already have equal denominators. Add the numerators and you MIGHT be done.